Circular No. 5937

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

GRB 940217 R. M. Kippen, J. Macri and J. Ryan, University of New Hampshire;

G. J. Fishman and C. Meegan, Space Sciences Laboratory, NASA/Marshall Space Flight Center; B. McNamara, New Mexico State University; V. Schoenfelder, Max-Planck Institut für Extraterrestrische Physik; W. Hermsen, SRON Laboratory for Space Research, Leiden; and K. Bennett, Astrophysics Division, Space Science Department, ESTEC, European Space Agency, report: “As part of the BATSE/COMPTEL/NMSU Rapid Burst Response Campaign, COMPTEL has imaged a strong cosmic γ-ray burst that was first detected by the BATSE experiment on Feb. 17.96021 UT. This event is one of the strongest bursts yet observed by BATSE and COMPTEL, with several hundred COMPTEL telescope events. Significant emission > 0.72 MeV was measured for > 150 s. Within this interval an intense, multipeaked structure is observed. The preliminary COMP-TEL imaging analysis yields a roughly circular uncertainty region centered at α = 1^{h}59^{m}.8, δ = +3^{o}23′ (equinox 2000.0; 3σ-confidence radius 1^{o}.2). Four hours after the burst the COMPTEL location was distributed to a worldwide network of multiwavelength observatories to search for fading counterparts. We encourage further follow-up observations as quickly as possible. Interested parties should contact B. McNamara at New Mexico State University (e-mail bmcnamar@nmsu.edu, telephone 505-646-2614).”

PERIODIC COMET SHOEMAKER-LEVY 9 (1993e) Measurements by R. H. McNaught from an image obtained by G. Da

Costa, F. Freeman, C. Heisler and T. Hill with the Anglo-Australian Telescope on Feb. 14 confirm all 21 nuclei in the Hubble Space Telescope images of Jan. 24–27 (IAUC 5936) and permit now the determination of an orbit for nucleus 3. This is given with new orbits for nuclei 2 and 4 on MPEC 1994-D07. The Jupiter-collision time for nucleus 3 is July 21.9 UT. The latest orbits (from MPC 23105–23107, MPEC 1994-D02 and D07) for 19 nuclei, together with daily ephemerides Feb. 17–July 17 (to July 16 for nucleus 21), are given in the Computer Service, option “X”.

J. J. Eitter, P. N. Appleton and R. J. Lavery, Iowa State University, report that the comet has faded to a low average surface brightness of 24.2 mag arcsec^{−2}. R-band CCD observations made at the Fick Observatory on Feb. 11.5 UT in conditions of good seeing show the six brightest nuclei to have faded to a surface brightness of 23.6–24.5 mag arcsec^{−2}. The comet’s integrated R-band magnitude was 15.4.

1994 February 21 Brian G. Marsden